OpenStudy (anonymous):

divide 142.20 in the ratio 1/4 : 1/5

OpenStudy (anonymous):

@mathman123 @hartnn

OpenStudy (anonymous):

@Jack1

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

79 and 63.2

OpenStudy (anonymous):

1/4:1/5 implies 1/4/(1/4+1/5)=5/9 and 1/5/(1/4+1/5)=4/9 142.2*5/9= 79 nad 142.2*4/9= 63.2

OpenStudy (anonymous):

I get 113.76 as the answer, is it right or wrong? here is my method: 1/4 : 1/5 = 0.25/0.2 = 1.25 or 5/4 142.20/1.25 = 113.76

OpenStudy (anonymous):

here we have to divide 142.2 in the proportion of x and y x being 1/4 and y being 1/5 we have to bring the denominator of x and y to same base. hence divide both x and y with (x+y) i.e., x/(x+y) and y/(x+y).Now multiply both with 142.2. this is ratio and proportion topic you can cross verify the solution as proportions should be added to 142.2

OpenStudy (anonymous):

ahha thanks minoz

OpenStudy (anonymous):

ahha thanks minoz

OpenStudy (anonymous):

welcome :)